# Statistical significance in choice modelling: computation, usage and reporting

### v2

@@ 4.3 Making decisions based on tests
A crucial step in the statistical test process is thus the decision the analyst makes for $\alpha_{\text{type I}}$, i.e. the selection of the significance level. It is very customary to set $\alpha_{\text{type I}}=0.05$, and many analysts use that level mechanically in their decisions of which parameters to retain in their models when looking at $t$-ratios, or to decide whether or not to reject a restricted model. Such a mechanical malpractice should be avoided. The selection of $\alpha_{\text{type I}}$ should depend on a mix of the plausibility of the null hypothesis and the cost of making a wrong decision. In classical hypothesis testing, the plausibility of the null hypothesis is very high, justifying small $\alpha_{\text{type I}}$, but this is not true for specification tests, such as the ones we face in model building.
@@
An additional point worth raising relates to the whether any corrections are needed to $p$-values as the estimation results of choice models imply the testing of multiple hypotheses at the same time (given that multiple parameters are estimated). In statistics, there is the notion that the probability of observing a rare event increases when conducting multiple tests, meaning that there would be increased risk of a type I error. Correction approaches have been put forward for this, including the Bonferroni correction (Bonferroni, 1936) and various false discovery rate (FDR) correction approaches (see e.g. Benjamini and Hochberg, 2018), with the view to increasing $p$-values in the case of multiple tests. Clearly, estimation results from a choice model relate to testing multiple hypotheses on a dataset, most obviously in terms of significance tests for each individual parameter, and if for example 20 parameters all obtain a $p$-value of 0.05, then one could expect that for one of these, the $H_{0}$ of no effect is rejected incorrectly. What specific correction, if any, should be applied is however unclear given that choice models rely on joint estimation of all effects. Any correction of the $p$-values would also increase the risk of type II errors and encourage analysts to incorrectly remove parameters from their models, putting them at risk of inferior model performance and endogeneity bias. Analysts should at the very least however stay clear of statements that all parameters are different from zero, given the above point.
@@ 4.4 Reporting of results
Remember that estimation gives values for $\hat{\beta_{k}}$ and $\hat{\sigma}_{k}$, i.e. estimates and standard errors. If an analyst includes these two outputs in the reporting of results, then this ensures that a reader is able to further process the results. The same applies if an analyst reports $\hat{\beta_{k}}$ and $t_{0,\hat{\beta}_{k}}$, i.e. the estimates and the $t$-ratios for parameters (against $\mathbf{H_{0}}$: $\beta_{k}=0$), as a reader can calculate $\hat{\sigma}_{k}=\frac{\hat{\beta_{k}}}{t_{0,\hat{\beta}_{k}}}$. From this perspective, reporting either standard errors or $t$-ratios raises no further issues in theory. Some analysts may prefer $t$-ratios over standard errors as it allows for more direct insights into whether the findings can reject $\mathbf{H_{0}}:\beta_{k}=0$ (and as they are dimension free), while others may prefer standard errors in order to compute confidence intervals or if they are interested in $t$-ratios against values other than 00.
[…]
@@
A key motivation for including $p$-values is clearly to allow readers to quickly see significance levels without having to compare $t$-ratios to critical values. The same applies to star measures, which are a somewhat more qualitative reporting tool as they simply split parameter estimates into four groups (i.e. $<90\%$ confidence, $90\%$ to $95\%$, $95\%$ to $99\%$, and $>99\%$). The use of these measures, accompanied by information on whether one-sided or two-sided tests are used, can thus be seen as providing additional information to readers. It is fine to report these measures alongside but not instead of standard errors or $t$-ratios, but this of course goes against the motivation often given for the use of star measures that it makes tables less cluttered and more readable.
@@ 5 Empirical example
The attributes of the alternatives used in the models for the present paper include in-vehicle travel time, out-of-vehicle travel time, and travel cost. We used a linear-in-attributes specification of a Multinomial Logit (MNL) model, with mode specific in vehicle and out of vehicle travel time parameters. All models were estimated using Apollo (Hess and Palma, 2019). In addition to the standard estimation (and computation of classical and robust standard errors), we also conducted bootstrap estimation using $400$ samples taken at the level of individual travellers (as opposed to individual trips).
[…]
@@
$\hat{\beta}_{k}$1.946 -0.4801 -4.927 -3.223 0.2193 -0.1236 -0.04388 -0.04328 -0.01418 -0.07446 -0.1107 -0.07922 -0.1088 -0.2230
$\hat{\sigma}_{k}$ 1.507E-01 3.119E-01 2.970E-01 2.140E-01 2.447E-01 9.631E-03 5.116E-03 9.736E-03 1.648E-02 8.783E-03 9.317E-03 9.391E-03 1.046E-02 1.638E-02
$\sigma_{robust,k}$ 4.977E-01 5.793E-01 8.850E-01 5.122E-01 4.431E-01 2.561E-02 9.346E-03 2.348E-02 5.087E-02 1.659E-02 2.002E-02 5.650E-02 2.129E-02 3.696E-02
$\sigma_{bootstrap,k}$ 4.772E-01 6.067E-01 1.014E 5.307E-01 4.451E-01 2.496E-02 9.365E-03 2.475E-02 6.181E-02 1.694E-02 2.076E-02 5.429E-02 2.274E-02 3.938E-02
[…]
$p_{t_{0,\beta_{k}},classical,1-sided}$ <.$<.001$** 0.062 * <.$<.001$** <.$<.001$** 0.185 <.$<.001$** <.$<.001$** <.$<.001$** 0.195 <.$<.001$** <.$<.001$** <.$<.001$** <.$<.001$** <.$<.001$**
$p_{t_{0,\beta_{k}},classical,2-sided}$ <.$<.001$** 0.1240 <.$<.001$** <.$<.001$** 0.370 <.$<.001$** <.$<.001$** <.$<.001$** 0.390 <.$<.001$** <.$<.001$** <.$<.001$** <.$<.001$** <.$<.001$**
[…]
$p_{t_{0,\beta_{k}},robust,1-sided}$ <.$<.001$** 0.2040 <.$<.001$** <.$<.001$** 0.310 <.$<.001$** <.$<.001$** 0.033 ** 0.390 <.$<.001$** <.$<.001$** 0.080 * <.$<.001$** <.$<.001$**
$p_{t_{0,\beta_{k}},robust,2-sided}$ <.$<.001$** 0.4070 <.$<.001$** <.$<.001$** 0.621 <.$<.001$** <.$<.001$** 0.065 * 0.780 <.$<.001$** <.$<.001$** 0.161 <.$<.001$** <.$<.001$**
[…]
$p_{t_{0,\beta_{k}},bootstrap,asymptotic1-sided}$ <.$<.001$** 0.2140 <.$<.001$** <.$<.001$** 0.311 <.$<.001$** <.$<.001$** 0.040 ** 0.409 <.$<.001$** <.$<.001$** 0.072 * <.$<.001$** <.$<.001$**
$p_{t_{0,\beta_{k}},bootstrap,asymptotic2-sided}$ <.$<.001$** 0.4290 <.$<.001$** <.$<.001$** 0.622 <.$<.001$** <.$<.001$** 0.080 * 0.819 <.$<.001$** <.$<.001$** 0.144 <.$<.001$** <.$<.001$**
$p_{t_{0,\beta_{k}},bootstrap,empirical1-sided}$ <.$<.001$** 0.2320 <.$<.001$** <.$<.001$** 0.263 <.$<.001$** <.$<.001$** 0.043 ** 0.374 <.$<.001$** <.$<.001$** 0.020 ** <.$<.001$** <.$<.001$**
$p_{LR}$ <.$<.001$** 0.1260 <.$<.001$** <.$<.001$** 0.368 <.$<.001$** <.$<.001$** <.$<.001$** 0.384 <.$<.001$** <.$<.001$** <.$<.001$** <.$<.001$** <.$<.001$**
$p_{LM}$ <.$<.001$** 0.1240 <.$<.001$** <.$<.001$** 0.370 <.$<.001$** <.$<.001$** <.$<.001$** 0.389 <.$<.001$** <.$<.001$** <.$<.001$** <.$<.001$** <.$<.001$**
$L_{\beta_{k},classical}$2.242E+00 -1.092E+00 -5.509E+00 -3.642E+00 -2.602E-01 -1.425E-01 -5.391E-02 -6.237E-02 -4.647E-02 -9.168E-02 -1.290E-01 -9.763E-02 -1.293E-01 -2.551E-01
$U_{\beta_{k},classical}$1.651E+00 1.313E-01 -4.345E+00 -2.803E+00 6.988E-01 -1.047E-01 -3.386E-02 -2.420E-02 1.812E-02 -5.725E-02 -9.244E-02 -6.082E-02 -8.833E-02 -1.909E-01
$L_{\beta_{k},robust}$2.922E+00 -1.616E+00 -6.662E+00 -4.226E+00 -6.491E-01 -1.738E-01 -6.220E-02 -8.931E-02 -1.139E-01 -1.070E-01 -1.499E-01 -1.900E-01 -1.506E-01 -2.955E-01
$U_{\beta_{k},robust}$9.709E-01 6.553E-01 -3.192E+00 -2.219E+00 1.088E+00 -7.339E-02 -2.557E-02 2.739E-03 8.553E-02 -4.194E-02 -7.145E-02 3.151E-02 -6.712E-02 -1.506E-01
$L_{\beta_{k},bootstrap}$2.793E+00 -1.672E+00 -7.$0$4.332E+00 -6.158E-01 -1.774E-01 -6.619E-02 -9.559E-02 -1.606E-01 -1.141E-01 -1.578E-01 -1.800E-01 -1.554E-01 -3.129E-01
$U_{\beta_{k},bootstrap}$9.863E-01 7.287E-01 -3.043E+00 -2.276E+00 1.141E+00 -8.045E-02 -2.847E-02 4.217E-03 7.661E-02 -4.570E-02 -7.413E-02 -5.520E-04 -7.422E-02 -1.552E-01
$L_{\beta_{k},HPD}$2.747E+00 -1.488E+00 -6.891E+00 -4.212E+00 -5.746E-01 -1.689E-01 -6.387E-02 -9.161E-02 -1.464E-01 -1.091E-01 -1.608E-01 -1.744E-01 -1.577E-01 -3.040E-01
$U_{\beta_{k},HPD}$9.574E-01 8.339E-01 -3.024E+00 -2.209E+00 1.150E+00 -7.474E-02 -2.653E-02 8.041E-03 7.949E-02 -4.120E-02 -7.887E-02 1.767E-03 -7.482E-02 -1.501E-01
@@
We first focus on the standard errors. For all $14$ model parameters, the classical standard errors ($\hat{\sigma}_{k}$) are substantially smaller than the robust standard errors ($\sigma_{robust,k}$), which is in line with general findings. The bootstrap standard errors ($\sigma_{bootstrap,k}$) are closer to the robust standard errors, which they exceed for $10$ out of $14$ parameters.
@@
We finally return to the issue raised early on about asymptotic normality. The use of the formula in Equation (10) uses the more stringent property of normality, which may not be appropriate outside the immediate neighbourhood of the optimum. To investigate this issue, we contrast the asymptotic confidence intervals with those obtained using bootstrapping. Of course, the bootstrap and robust confidence intervals are wider than those obtained using the classical standard errors, in line with the difference in the standard errors. However, we also note that the bootstrap confidence intervals are not necessarily symmetric around the MLE. We calculate an asymmetry index as $\frac{(U-M)-(M-L)}{U-L}$, where 00 would indicate symmetry. For the ASCs, we see negative skew for the CIs for $\delta_{taxi}$ and $\delta_{cycling}$, with positive skew for the remaining constants. More importantly, the CIs are negatively skewed for all $9$ parameters relating to travel time and travel costs. Comparing the empirical bootstrap CIs with the HDP intervals, we see that the latter are narrower for $13$ out of $14$ parameters, in line with earlier discussions. A majority of the CIs retain a negative skew, but this now changes to positive for $\delta_{cycling}$, $\beta_{car,tt}$ and $\beta_{rail,tt}$.

### v3

@@ 4.6 Making decisions based on tests
As discussed earlier, statistical tests are used to decide whether there is enough evidence in support of $\mathbf{H_{1}}$ to reject $\mathbf{H_{1}}$. In this process, an analyst needs to balance the risks of type I and type II errors. A crucial step in the statistical test process is the decision the analyst makes for $\alpha_{\text{type I}}$, i.e. the selection of the significance level. It is very customary to set $\alpha_{\text{type I}}=0.05$, and many analysts use that level mechanically in their decisions of which parameters to retain in their models when looking at $t$-ratios, or to decide whether or not to reject a restricted model. Such a mechanical malpractice should be avoided. The selection of $\alpha_{\text{type I}}$ should depend on a mix of the plausibility of the null hypothesis and the cost of making a wrong decision. In classical hypothesis testing, the plausibility of the null hypothesis is very high, justifying small $\alpha_{\text{type I}}$, but this is not true for specification tests, such as the ones we face in model building.
@@
The point about $\alpha_{\text{type I}}$ being the significance of a test also highlights a common bad practice. Analysts routinely describe the findings of statistical tests as showing that a parameter is “significantly different from 0”. It should first be noted that hypothesis testing is only concerned with whether the null hypothesis can be rejected, not whether the alternative hypothesis is true. In other words, it establishes whether our finding could be due to chance and that the true value is in fact zero (if $\mathbf{H_{0}}$: $\beta_{k}=0$). What is commonly described as a test of significance is thus a test to see whether the null hypothesis of a parameter being equal to zero can be rejected. The statistical significance of the test relates to the probability of $H_{0}$ being rejected when it is in fact true, i.e. a type I error. It is thus incorrect to talk about a parameter being $95\%$ significant. The significance level in this case is $5\%$, and we can reject $H_{0}$ with a $95\%$ level of confidence, not significance. The relationship with the earlier discussion on confidence intervals is clear. Using a negative estimate of $\hat{\beta}$ as our example, if we can reject $H_{0}:\beta=0$ with a confidence level of $C\%$ for a two-sided test, then this means that 0 forms the upper bound of the $C\%$ confidence interval around the estimated value $\hat{\beta_{k}}$.
It is also important to note that the definition of $\alpha_{\text{type I}}$ depends on the parametric sampling distribution, which varies by the type of test, holds asymptotically for DCM, and relies on specific assumptions. Consequently, the true (empirical) $\alpha_{\text{type I}}$ may differ from the calculated $\alpha_{\text{type I}}$. This difference is referred to as the size distortion of the statistical test. If the distortion is positive or “liberal”, the empirical $\alpha_{\text{type I}}$ is larger than the calculated one, and this increases the likelihood of committing more Type I errors than expected. Conversely, if the distortion is negative or “conservative”, the test makes fewer Type I errors than planned, but becomes more prone to Type II errors. Size distortion can be analyzed using Monte Carlo simulations, which replicate the resampling process and provide a useful tool for comparing the performance of different statistical tests.
A different perspective on the problem lies in considering the scientific value that emerges from the outcome of a test, the impact of significance on beliefs. Although it is usual practice to confer point rejections of the null (typically a p-value above 5%) a larger scientific value, Abadie (2020) remarks that the opposite is true instead. Given that often the prior of a rejection is larger than 50%, a non-significant result (i.e. accepting the null) is more surprising, and consequently more informative than a significant result.
@@
An additional point worth raising relates to the whether any corrections are needed to $p$-values as the estimation results of choice models imply the testing of multiple hypotheses at the same time (given that multiple parameters are estimated). In statistics, there is the notion that the probability of observing a rare event increases when conducting multiple tests, meaning that there would be increased risk of a type I error. Correction approaches have been put forward for this, including the Bonferroni correction (Bonferroni, 1936) and various false discovery rate (FDR) correction approaches (see e.g. Benjamini and Hochberg, 2018), with the view to increasing $p$-values in the case of multiple tests. Clearly, estimation results from a choice model relate to testing multiple hypotheses on a dataset, most obviously in terms of significance tests for each individual parameter, and if for example 20 parameters all obtain a $p$-value of 0.05, then one could expect that for one of these, the $H_{0}$ of no effect is rejected incorrectly. What specific correction, if any, should be applied is however unclear given that choice models rely on joint estimation of all effects. Any correction of the $p$-values would also increase the risk of type II errors and encourage analysts to incorrectly remove parameters from their models, putting them at risk of inferior model performance and endogeneity bias. Analysts should at the very least however stay clear of statements that all parameters are different from zero, given the above point. For this reason, joint hypothesis tests, such as likelihood ratio tests, are generally better suited for assessing the relevance of groups of parameters in choice models than multiple individual $t$-tests, as they typically exhibit greater power for joint hypotheses and do not rely on ad-hoc multiple-testing corrections.
@@ 4.7 Reporting of results
Remember that estimation gives values for $\hat{\beta_{k}}$ and $\hat{\sigma}_{k}$, i.e. estimates and standard errors. If an analyst includes these two outputs in the reporting of results, then this ensures that a reader is able to further process the results. The same applies if an analyst reports $\hat{\beta_{k}}$ and $t_{0,\hat{\beta}_{k}}$, i.e. the estimates and the $t$-ratios for parameters (against $\mathbf{H_{0}}$: $\beta_{k}=0$), as a reader can calculate $\hat{\sigma}_{k}=\frac{\hat{\beta_{k}}}{t_{0,\hat{\beta}_{k}}}$. From this perspective, reporting either standard errors or $t$-ratios raises no further issues in theory. Some analysts may prefer $t$-ratios over standard errors as it allows for more direct insights into whether the findings can reject $\mathbf{H_{0}}:\beta_{k}=0$ (and as they are dimension free), while others may prefer standard errors in order to compute confidence intervals or if they are interested in $t$-ratios against values other than 0.
A practical consideration arises in terms of the numerical precision that is used in presenting the results. As standard errors are often an order of magnitude smaller than parameter estimates, the use of too low a number of digits will mean valuable information is lost in reporting results. Analysts should ensure to report at least two significant digits (i.e. not counting leading zeros). For $t$-ratios, it is common to report two digits, although the second decimal might be considered less useful as it relates to 1% of a standard error. For parameter estimates, analysts may choose to report even more digits for special cases, while realising that reporting too large a number of digits claims spurious accuracy and confuses the reader.
The situation with reporting becomes more complicated if an analyst reports $p$-values instead of $t$-ratios and standard errors, as is common practice in some disciplines, such as health, while being rare in others, such as transport. First, given the above points about one-sided vs two-sided tests, if an analyst reports $p$-values only (i.e. without standard errors and/or $t$-ratios), then, in order to avoid misinterpretation, information must be provided on whether these relate to a one-sided or a two-sided test. Second, an issue with precision and further use of the results arises. Let us assume that we have $\mid t_{0,\hat{\beta}}\mid>10$, in which case, whether using a one-sided or a two-sided test, an analyst is likely going to report $p=0$ or $p<10^{-5}$ or, preferably, $p<.001$, the APA style option. Of course, the test clearly rejects $H_{0}:\beta=0$, but reporting $p$-values alone will prevent a reader from using the results to compute confidence intervals or to gain any insights into differences in uncertainty for two parameters which both have $p<10^{-5}$.
@@
A key motivation for including $p$-values is clearly to allow readers to quickly see significance levels without having to compare $t$-ratios to critical values. The same applies to star measures, which are a somewhat more qualitative reporting tool as they simply split parameter estimates into four groups (i.e. $90\%$ confidence, $90\%$ to $\%, 9595\%$ to $99\%$, and $99\%$). The use of these measures, accompanied by information on whether one-sided or two-sided tests are used, can thus be seen as providing additional information to readers. It is fine to report these measures alongside but not instead of standard errors or $t$-ratios, but this of course goes against the motivation often given for the use of star measures that it makes tables less cluttered and more readable.
@@ 5 Empirical example
The attributes of the alternatives used in the models for the present paper include in-vehicle travel time, out-of-vehicle travel time, and travel cost. We used a linear-in-attributes specification of a Multinomial Logit (MNL) model, with mode specific in vehicle and out of vehicle travel time parameters. For this illustration, we rely on such a simple model, including without an explicit treatment of the repeated choice nature of the data. All models were estimated using Apollo (Hess and Palma, 2019) [^14] The data and code are available via www.stephanehess.me.uk/papers/significance.zip.. In addition to the standard estimation (and computation of classical and robust standard errors), we also conducted bootstrap estimation using $400$ samples taken at the level of individual travellers (as opposed to individual trips).
The results of our analysis are shown in Table 3. We follow the policy of showing at least 2 significant digits (i.e. not counting leading zeros) for estimates, standard errors and confidence limits, which in our case is possible by showing four digits after the decimal point throughout, and 5 in one case. For the latter two, we use scientific notation giving the larger number of leading zeros for some parameters. For p-values, we show three digits after the decimal point, adopting the APA convention of $p<.001$ where this applies.
@@
$\hat{\beta}_{k}$ -1.9464 -0.4801 -4.9270 -3.2225 0.2193 -0.1236 -0.0439 -0.0433 -0.0142 -0.0745 -0.1107 -0.0792 -0.1088 -0.2230
$\hat{\sigma}_{k}$ 0.1507 0.3119 0.2970 0.2140 0.2447 0.0096 0.0051 0.0097 0.0165 0.0088 0.0093 0.0094 0.0105 0.0164
$\sigma_{robust,k}$ 0.4977 0.5793 0.8850 0.5122 0.4431 0.0256 0.0093 0.0235 0.0509 0.0166 0.0200 0.0565 0.0213 0.0370
$\sigma_{bootstrap,k}$ 0.4772 0.6067 1.0139 0.5307 0.4451 0.0250 0.0094 0.0248 0.0618 0.0169 0.0208 0.0543 0.0227 0.0394
$t_{0,\beta_{k}}\text{ (classical)}$ -12.91 -1.54 -16.59 -15.06 0.90 -12.83 -8.58 -4.45 -0.86 -8.48 -11.88 -8.44 -10.40 -13.61
$p_{t_{0,\beta_{k}},classical,1-sided}$ <.001<.001 *** 0.062 * <.001<.001 *** <.001<.001 *** 0.185 <.001<.001 *** <.001<.001 *** <.001<.001 *** 0.195 <.001<.001 *** <.001<.001 *** <.001<.001 *** <.001<.001 *** <.001<.001 ***
$p_{t_{0,\beta_{k}},classical,2-sided}$ <.001<.001 *** 0.124 <.001<.001 *** <.001<.001 *** 0.370 <.001<.001 *** <.001<.001 *** <.001<.001 *** 0.390 <.001<.001 *** <.001<.001 *** <.001<.001 *** <.001<.001 *** <.001<.001 ***
$t_{0,\beta_{k}}\text{ (robust)}$ -3.91 -0.83 -5.57 -6.29 0.49 -4.83 -4.70 -1.84 -0.28 -4.49 -5.53 -1.40 -5.11 -6.03
$p_{t_{0,\beta_{k}},robust,1-sided}$ <.001<.001 *** 0.204 <.001<.001 *** <.001<.001 *** 0.310 <.001<.001 *** <.001<.001 *** 0.033 ** 0.390 <.001<.001 *** <.001<.001 *** 0.080 * <.001<.001 *** <.001<.001 ***
$p_{t_{0,\beta_{k}},robust,2-sided}$ <.001<.001 *** 0.407 <.001<.001 *** <.001<.001 *** 0.621 <.001<.001 *** <.001<.001 *** 0.065 * 0.780 <.001<.001 *** <.001<.001 *** 0.161 <.001<.001 *** <.001<.001 ***
$t_{0,\beta_{k}}\text{ (bootstrap)}$ -4.08 -0.79 -4.86 -6.07 0.49 -4.95 -4.69 -1.75 -0.23 -4.40 -5.33 -1.46 -4.79 -5.66
$p_{t_{0,\beta_{k}},bootstrap,asymptotic1-sided}$ <.001<.001 *** 0.214 <.001<.001 *** <.001<.001 *** 0.311 <.001<.001 *** <.001<.001 *** 0.040 ** 0.409 <.001<.001 *** <.001<.001 *** 0.072 * <.001<.001 *** <.001<.001 ***
$p_{t_{0,\beta_{k}},bootstrap,asymptotic2-sided}$ <.001<.001 *** 0.429 <.001<.001 *** <.001<.001 *** 0.622 <.001<.001 *** <.001<.001 *** 0.080 * 0.819 <.001<.001 *** <.001<.001 *** 0.144 <.001<.001 *** <.001<.001 ***
$p_{t_{0,\beta_{k}},bootstrap,empirical1-sided}$ <.001<.001 *** 0.232 <.001<.001 *** <.001<.001 *** 0.263 <.001<.001 *** <.001<.001 *** 0.043 ** 0.374 <.001<.001 *** <.001<.001 *** 0.020 ** <.001<.001 *** <.001<.001 ***
$p_{LR}$ <.001<.001 *** 0.126 <.001<.001 *** <.001<.001 *** 0.368 <.001<.001 *** <.001<.001 *** <.001<.001 *** 0.384 <.001<.001 *** <.001<.001 *** <.001<.001 *** <.001<.001 *** <.001<.001 ***
$p_{LM}$ <.001<.001 *** 0.124 <.001<.001 *** <.001<.001 *** 0.370 <.001<.001 *** <.001<.001 *** <.001<.001 *** 0.389 <.001<.001 *** <.001<.001 *** <.001<.001 *** <.001<.001 *** <.001<.001 ***
$L_{\beta_{k},classical}$ -2.2419 -1.0915 -5.5092 -3.6418 -0.2602 -0.1425 -0.0539 -0.0624 -0.0465 -0.0917 -0.1290 -0.0976 -0.1293 -0.2551
$U_{\beta_{k},classical}$ -1.6510 0.1313 -4.3448 -2.8032 0.6988 -0.1047 -0.0339 -0.0242 0.0181 -0.0572 -0.0924 -0.0608 -0.0883 -0.1909
$L_{\beta_{k},robust}$ -2.9219 -1.6156 -6.6616 -4.2264 -0.6491 -0.1738 -0.0622 -0.0893 -0.1139 -0.1070 -0.1499 -0.1900 -0.1506 -0.2955
$U_{\beta_{k},robust}$ -0.9709 0.6553 -3.1925 -2.2186 1.0877 -0.0734 -0.0256 0.0027 0.0855 -0.0419 -0.0715 0.0315 -0.0671 -0.1506
$L_{\beta_{k},bootstrap}$ -2.7926 -1.6724 -7.0079 -4.3318 -0.6158 -0.1774 -0.0662 -0.0956 -0.1606 -0.1141 -0.1578 -0.1800 -0.1554 -0.3129
$U_{\beta_{k},bootstrap}$ -0.9863 0.7287 -3.0426 -2.2758 1.1413 -0.0804 -0.0285 0.0042 0.0766 -0.0457 -0.0741 -0.00055 -0.0742 -0.1552
$L_{\beta_{k},HPD}$ -2.7472 -1.4885 -6.8906 -4.2116 -0.5746 -0.1689 -0.0639 -0.0916 -0.1464 -0.1091 -0.1608 -0.1744 -0.1577 -0.3040
$U_{\beta_{k},HPD}$ -0.9574 0.8339 -3.0243 -2.2092 1.1497 -0.0747 -0.0265 0.0080 0.0795 -0.0412 -0.0789 0.0018 -0.0748 -0.1501
@@
$^{***}:p\leq 0.01$; $^{**}:p\leq 0.05$; $^{*}:p\leq 0.1$
We first focus on the standard errors. For all $14$ model parameters, the classical standard errors ($\hat{\sigma}_{k}$) are substantially smaller than the robust standard errors ($\sigma_{robust,k}$), which is in line with general findings, and is likely in part related to the lack of an explicit treatment of the repeated choice nature of the data in our model. The bootstrap standard errors ($\sigma_{bootstrap,k}$) are closer to the robust standard errors, which they exceed for $10$ out of $14$ parameters.
@@
We finally return to the issue raised early on about asymptotic normality. The use of the formula in Equation (11) uses the more stringent property of normality, which may not be appropriate outside the immediate neighbourhood of the optimum. To investigate this issue, we contrast the asymptotic confidence intervals with those obtained using bootstrapping. Of course, the bootstrap and robust confidence intervals are wider than those obtained using the classical standard errors, in line with the difference in the standard errors. However, we also note that the bootstrap confidence intervals are not necessarily symmetric around the MLE. We calculate an asymmetry index as $\frac{(U-M)-(M-L)}{U-L}$, where 0 would indicate symmetry. For the ASCs, we see negative skew for the CIs for $\delta_{taxi}$ and $\delta_{cycling}$, with positive skew for the remaining constants. More importantly, the CIs are negatively skewed for all $9$ parameters relating to travel time and travel costs. Comparing the empirical bootstrap CIs with the HDP intervals, we see that the latter are narrower for $13$ out of $14$ parameters, in line with earlier discussions. A majority of the CIs retain a negative skew, but this now changes to positive for $\delta_{cycling}$, $\beta_{car,tt}$ and $\beta_{rail,tt}$.